# Language-centered Human Activity Recognition

### v2

@@ Abstract.
Human Activity Recognition (HAR) using Inertial Measurement Unit (IMU) sensors is critical for applications in healthcare, safety, and industrial production. However, variations in activity patterns, device types, and sensor placements create distribution gaps across datasets, reducing the performance of HAR models. To address this, we propose LanHAR, a novel system that leverages Large Language Models (LLMs) to generate semantic interpretations of sensor readings and activity labels for cross-dataset HAR. This approach not only mitigates cross-dataset heterogeneity but also enhances the recognition of new activities. LanHAR employs an iterative re-generation method to produce high-quality semantic interpretations with LLMs and a two-stage training framework that bridges the semantic interpretations of sensor readings and activity labels. This ultimately leads to a lightweight sensor encoder suitable for mobile deployment, enabling any sensor reading to be mapped into the semantic interpretation space. Experiments on four public datasets demonstrate that our approach significantly outperforms state-of-the-art methods in both cross-dataset HAR and new activity recognition. The source code will be made publicly available.
Human activity recognition, Natural language, Large language models
@@ 1. Introduction
Existing work on cross-dataset HAR can be broadly categorized into two approaches: domain adaptation and data augmentation. Domain adaptation methods (Chang et al., 2020; Qin et al., 2019; Khan et al., 2018; Zhou et al., 2020) aim to bridge the distribution gap between source and target datasets to enhance model performance in training. Data augmentation methods (Um et al., 2017; Xu et al., 2023; Qian et al., 2022; Saeed et al., 2019) seek to increase the diversity of training data to improve the model’s generalization ability. While both approaches are valuable, they still suffer from two limitations: (i) Most existing work overlooks the physical meanings of activities. This can exacerbate the impact of classification errors on downstream applications. For instance, misclassifying a walking activity as running may have a relatively minor effect compared to misclassifying it as a significantly different activity such as biking (see quantitative results in Evaluation). (ii) Moreover, cross-dataset scenarios often introduce new activities that these methods struggle to handle, as these new activities are typically missing from the training phase.
@@
We argue that semantic interpretations of sensor readings and activity labels can play a crucial role in addressing the aforementioned limitations as shown in Figure 1. (i) For cross-dataset heterogeneity, interpreting sensor readings based on its physical meanings can help identify patterns associated with specific activities (e.g., jogging often exhibits regular periodicity). The underlying intuition is that, despite the variability across datasets, activities may still reveal common underlying patterns. By identifying these patterns, we can map the heterogeneous data into a common space, thereby mitigating the impact of dataset variability. (ii) When dealing with new activities, semantic interpretations of activity labels can enhance their meaning by generating detailed descriptions. Combining semantic interpretations of both sensor readings and activity labels creates the potential to map new data to new activity labels. However, manual semantic interpretation is resource-intensive. The recent success of large language models (LLMs) offers a promising solution, enabling the generation of semantic interpretations from both sensor readings and activity labels through natural language (Jiang et al., 2024; Zhang et al., 2023). Research has shown that LLMs have the ability to perceive aspects of the physical world (Xu et al., 2024), providing an opportunity to leverage these capabilities to bridge the gap in cross-dataset HAR through automated semantic interpretation.
@@
Despite the promise of semantic interpretations using LLMs, several challenges remain to be addressed: (i) Generation Challenge: The semantic interpretations generated by LLMs can suffer from issues like hallucinations and randomness. Ensuring consistent and high-quality semantic interpretations remains a significant challenge. (ii) Alignment Challenge: Once LLMs generate semantic interpretations of sensor readings and labels, the next step is to align them for Human Activity Recognition (HAR), i.e., aligning sensor reading and label based on their interpretation. While language models can be used to align these interpretations, general LLMs lack specific knowledge of IMU data and human activity recognition, limiting their ability to accurately understand and reason about activities. (iii) Deployment Challenge: Frequent access to cloud-based LLMs or deploying them locally on mobile devices is impractical due to high latency and resource constraints. A lightweight framework is needed to effectively leverage LLMs for on-device HAR.
@@
To address these challenges, we design LanHAR  a novel system that leverages LLMs to generate semantic interpretations of sensor readings and activity labels, aimed at improving cross-dataset human activity recognition. LanHAR consists of three key components. (i) Generation: we carefully design prompts to guide LLMs in generating semantic interpretations and develop an iterative re-generation method to ensure high-quality semantic outputs. (ii) Alignment: we design a semantic interpretation alignment module to enhance understanding of semantic interpretations based on a pre-trained text encoder to encode interpretations and two contrastive learning tasks to improve the alignment between activities and sensor readings. (iii) Deployment: We introduce a lightweight, two-stage training and inference framework. First, we train the semantic interpretation alignment module, then design and train a sensor reading encoder to map sensor data to the semantic interpretation space. For inference on mobile devices, we use the trained encoder to obtain sensor reading encodings and measure their similarity to activity label semantics to determine the HAR results.
@@
  • To our knowledge, we are the first to leverage LLMs to generate semantic interpretations for cross-dataset human activity recognition. Our approach introduces a novel perspective by utilizing semantic interpretations from both sensor readings and activity labels, which helps mitigate the impact of cross-dataset heterogeneity and opens the door to recognizing new activities.
  • We design LanHAR, a novel system that leverages LLMs to generate semantic interpretations of sensor readings and activity labels, addressing the challenge of cross-dataset human activity recognition. The system features a semantic interpretation generation process with an iterative re-generation method to ensure high-quality outputs, alongside a two-stage training framework that transfers the capabilities of large language models (LLMs) to mobile devices.
  • We evaluate our system based on four public datasets. The experimental results demonstrate that it outperforms other state-of-the-art methods in cross-dataset activity recognition, achieving an improvement of 7.21% in accuracy and 13.31% in F1 score. Additionally, for new activity recognition, it shows further enhancement, with improvements of 74.65% in accuracy.
@@ 2.1. LLMs possess the ability to perceive the physical world
Existing research (Xu et al., 2024; Ji et al., 2024) has demonstrated that LLMs possess the ability to perceive and interpret the physical world. For instance, LLMs can analyze various types of sensor data to infer a person’s location or activity. Building on this capability, we explore the use of LLMs to analyze sensor readings and describe activity labels, thereby generating semantic interpretations for both. Specifically, we design prompts that include data introduction, data analysis, relevant knowledge, and task instructions (detailed in Section 3.3) to guide LLMs in generating semantic interpretations of sensor readings. Similarly, we guide LLMs to produce descriptions related to activity labels. Figure 2 provides an example where GPT-4 generates its understanding of jogging alongside an analysis of sensor data sequences.
@@ 2.2. Semantic interpretations of sensor readings for cross-dataset HAR
We explain why LLM-generated semantic interpretation offers a valuable approach to addressing data heterogeneity in cross-dataset HAR. The key intuition is that, despite variations in how heterogeneous data manifests, it often has underlying common activity patterns. By identifying these shared patterns, we can map the original heterogeneous sensor data into a common shared space (i.e., a language space with encoded semantic interpretation), effectively reducing the impact of data heterogeneity.
@@
To understand how data heterogeneity is mitigated using generated semantic interpretations, we compare data distributions under three settings: raw data, data encoded by encoders, and generated semantic interpretations. Data encoding is a common approach used in existing research (Auge et al., 2021), aimed at learning better representations or embeddings of sensor readings. For data encoding, we employ a self-supervised learning model, BERT (Devlin et al., 2018), as a representative method, following prior work (Xu et al., 2021) that masks and predicts parts of the sensor readings. For semantic interpretation, we use GPT-4 to generate interpretations from raw IMU data based on a designed prompt (detailed in Section 3.3). These interpretations are then input into a pre-trained language model (BERT) to produce representations in the language space. We use two public datasets—UCI (Reyes-Ortiz et al., 2016) (D1) and Motion (Malekzadeh et al., 2019) (D2) as a cross-dataset example including four common activities: walking, sitting, going upstairs, and going downstairs.
@@
Qualitative results: We visualize the results from the three settings using t-SNE to project them onto a 2D surface. As an example, we plot two activities, “sit” (blue) and “walk” (orange), in Figure 3. Initially, in the raw data, the same activities from different datasets are far apart. After applying self-supervised learning, the distribution gap between the datasets decreases, but the activities also become less distinguishable. However, when using semantic interpretations, the distribution gap reduces further, and the same activities (same color) cluster more closely, demonstrating improved grouping of the same activities.
@@
Quantitative results: We quantitatively assess the impact of introducing semantic interpretations for sensor readings across all activities. To measure the distribution differences between the two datasets under the three settings, we use Kullback-Leibler (KL) divergence, a widely used metric for comparing distributions. Table 1 presents the KL divergence for all activities across the two datasets under each setting. The results show that introducing semantic interpretations significantly reduces the distribution gap for all activities. On average, the KL divergence is reduced by 56.89%.
@@
Table 1. KL divergence between the distributions of the same activity across two datasets
@@
Walking 0.0194 0.0066 0.0013
Sitting 0.0478 0.0319 0.0102
Going upstairs 0.0260 0.0118 0.0062
Going downstairs 0.0303 0.0163 0.0112
Average 0.0309 0.0167 0.0072
@@ 2.3. New activity recognition
Recognizing new activities has been a longstanding challenge in HAR, particularly for machine learning-based methods, which are unable to identify activities that were not present during the training phase. The most relevant approach to address this challenge involves using predefined attributes of activities to establish connections between new and existing activities (Cheng et al., 2013). However, the performance of this method relies heavily on the quality of manually defined activity attributes, making it costly and difficult to generalize to new labels.
We argue that combining the semantic interpretations of sensor readings and labels offers a new perspective for addressing this challenge. By converting all activity labels into a language space using their semantic interpretations, we can represent labels even without corresponding sensor readings (which is common for new labels). Similarly, with sensor readings mapped into the same language space, we can match each reading to a label, regardless of whether the label was previously encountered. Figure 4 illustrates this concept, where we assume “biking” and “going upstairs” are known activities from the training phase, while “going downstairs” is a new activity. In our approach, we first convert all activity labels into the same language space using their LLM-generated semantic interpretations. We then obtain semantic interpretations for the sensor readings. Even though “going downstairs” is unseen during training, we can still measure its distance to all activity labels in the language space, enabling the recognition of new activities. In our setting, we assume that we have the knowledge of possible activity labels in the inference stage.
@@ 3.1. Problem formulation
Figure 5. Overview of LanHAR. (1) Utilize LLMs for generating semantic interpretations of sensor reading and activity labels. (2) Train a text encoder to encode two types of semantic interpretations and achieve their alignment. $H{i}$ and $Z_{i}$ denote embeddings of semantic interpretations of activity labels and sensor reading (3) Train a sensor encoder to align sensor reading and semantic interpretation. $E{i}$ denote embeddings of sensor reading (4) For inference, only use the sensor encoder to generate embeddings for the sensor readings $E{i}$ and then compute the similarity with the pre-stored embeddings of the activity labels $H{i}$ to obtain the human activity recognition results.
@@ 3.2. Design overview
We design LanHAR, a novel system that leverages LLM-generated semantic interpretations of sensor readings and activity labels for cross-dataset human activity recognition. As shown in Figure 5, LanHAR consists of four processes:
@@ 3.3. LLMs for semantic interpretations
We introduce how we guide LLMs to generate semantic interpretations of sensor readings and activity labels. We also discuss our approach to handling low-quality interpretations to ensure the accuracy and reliability of LLM responses.
@@ 3.3.1. Obtaining semantics interpretation of sensor reading
We carefully design a prompt to help LLMs better understand the problem setting and deliver more accurate and meaningful answers. An example is shown in Figure 6. The prompt consists of four key parts: data introduction, data analysis, knowledge, and task introduction.
@@ 3.3.2. Obtaining semantics interpretation of activity label
Similarly, we design a prompt to guide LLMs in generating semantic interpretations of activity labels. In this prompt, we outline the task for the LLMs, instructing them to generate a description for a given activity. The description should cover three key aspects: a general overview of the activity, the potential states or patterns detected by the accelerometer and gyroscope during the activity, and the body parts likely involved in performing the activity. An example of this prompt is shown in Figure 7.
@@ 3.3.3. An iterative re-generation method to ensuring the quality of LLM responses
The quality of LLMs’ interpretations is critical for our approach. However, due to hallucinations or other incidental factors, LLMs may sometimes generate inaccurate or illogical responses. Through manual review, we identified many such cases when generating semantic interpretations. To improve the accuracy of LLM responses, we developed an iterative re-generation method to ensure high-quality outputs. The key intuition is that filtering out low-quality responses will result in a lower KL divergence within the same activity. As illustrated in Figure 8, the framework first identifies the $k$ most problematic responses based on KL divergence. These data points are re-input into the LLM to regenerate the semantic interpretations. We then assess whether to incorporate the new interpretations based on their ability to reduce the overall KL divergence. The details are as follows. We show in the evaluation this process improves the data quality.
Figure 8. The iterative re-generatio method to ensuring the quality of LLM responses. (1) Filter inaccurate semantic interpretations. (2) Regenerate new semantic interpretations by LLMs. (3) Incorporate the new semantic interpretation based on whether it can reduce the overall KL divergence.
(1) Filter out inaccurate semantic interpretations: We randomly divide the generated semantic interpretations of the same activity into two parts, $A$ and $B$, and calculate the initial KL divergence between them. Next, we perform an iterative selection process: in each iteration, we identify and select the data sample from one part that most negatively impacts of KL divergence, adding it to a selection set. After selecting a data sample, it is removed from its original part. This process is repeated until the selection set contains $k$ data samples.
(2) Regenerate new semantic interpretations by LLMs: After identifying the $k$ data samples, we update the task introduction in the prompt (Section 3.3.1) to inform the LLMs of the inaccurate responses. The revised prompt includes the following: “This is your previous response to the task. Please analyze it step by step to identify any logical errors, inconsistencies with real-world knowledge, or discrepancies with the input data. Provide a corrected response according to the required format.”
(3) Incorporate the new semantic interpretation based on whether it can reduce the overall KL divergence: If the newly generated interpretations reduce the KL divergence, we accept the new semantic interpretation; otherwise, we retain the original one. We repeat this process until either the KL divergence no longer shows significant improvement for multiple continuous iterations or the specified number of iterations is reached.

### v3

@@ Abstract.
Human Activity Recognition (HAR) using Inertial Measurement Unit (IMU) sensors is critical for applications in healthcare, safety, and industrial production. However, variations in activity patterns, device types, and sensor placements create distribution gaps across datasets, reducing the performance of HAR models. To address this, we propose LanHAR, a novel system that leverages Large Language Models (LLMs) to generate semantic interpretations of sensor readings and activity labels for cross-dataset HAR. This approach not only mitigates cross-dataset heterogeneity but also enhances the recognition of new activities. LanHAR employs an iterative re-generation method to produce high-quality semantic interpretations with LLMs and a two-stage training framework that bridges the semantic interpretations of sensor readings and activity labels. This ultimately leads to a lightweight sensor encoder suitable for mobile deployment, enabling any sensor reading to be mapped into the semantic interpretation space. Experiments on five public datasets demonstrate that our approach significantly outperforms state-of-the-art methods in both cross-dataset HAR and new activity recognition. The source code will be made publicly available.
Human activity recognition, Natural language processing, Large language models
copyright: noneconference: ; ; ccs: Human-centered computing Ubiquitous and mobile computing
@@ 1. Introduction
Existing work on cross-dataset HAR can be broadly categorized into two methods: domain adaptation and data augmentation. Domain adaptation methods (Chang et al., 2020; Qin et al., 2019; Khan et al., 2018; Zhou et al., 2020) aim to bridge the distribution gap between source and target datasets to enhance model performance in training. Data augmentation methods (Um et al., 2017; Xu et al., 2023; Qian et al., 2022; Saeed et al., 2019) seek to increase the diversity of training data to improve the model’s generalization ability. While both methods are valuable, they still suffer from two limitations: (i) Most existing work overlooks the physical meanings of activities. This can exacerbate the impact of classification errors on downstream applications. For instance, misclassifying a walking activity as running may have a relatively minor effect compared to misclassifying it as a significantly different activity such as biking (see quantitative results in Evaluation). (ii) Moreover, cross-dataset scenarios often introduce new activities that these methods struggle to handle, as these new activities are typically missing from the training phase.
We argue that semantic interpretations of sensor readings and activity labels can play a crucial role in addressing the aforementioned limitations as shown in Fig. 1. (i) For cross-dataset heterogeneity, interpreting sensor readings based on its physical meanings can help identify patterns associated with specific activities (e.g., jogging often
@@
exhibits regular periodicity). The underlying intuition is that, despite the variability across datasets, activities may still reveal common underlying patterns. By identifying these patterns, we can map the heterogeneous data into a common space, thereby mitigating the impact of dataset variability. (ii) When dealing with new activities, semantic interpretations of activity labels can enhance their meaning by generating detailed descriptions. Combining semantic interpretations of both sensor readings and activity labels creates the potential to map new data to new activity labels. However, manual semantic interpretation is resource-intensive. The recent success of large language models (LLMs) offers a promising solution, enabling the generation of semantic interpretations from both sensor readings and activity labels through natural language (Jiang et al., 2024; Zhang et al., 2023). Research has shown that LLMs have the ability to perceive aspects of the physical world (Xu et al., 2024), providing an opportunity to leverage these capabilities to bridge the gap in cross-dataset HAR through automated semantic interpretation.
@@
Despite the promise of semantic interpretations using LLMs, several challenges remain to be addressed: (i) Generation Challenge: The semantic interpretations generated by LLMs can suffer from issues like hallucinations and randomness. Ensuring consistent and high-quality semantic interpretations remains a significant challenge. (ii) Alignment Challenge: Once LLMs generate semantic interpretations of sensor readings and labels, the next step is to align them for human activity recognition, i.e., aligning sensor readings and labels based on their interpretations. While language models can be used to align these interpretations, general language models lack specific knowledge of IMU data and human activity recognition, limiting their ability to accurately understand and reason about activities. (iii) Deployment Challenge: Frequent access to cloud-based LLMs or deploying them locally on mobile devices is impractical due to high latency and resource constraints. A lightweight framework is needed to effectively leverage LLMs for on-device HAR.
@@
To address these challenges, we design LanHAR  a novel system that leverages LLMs to generate semantic interpretations of sensor readings and activity labels, aiming to improve cross-dataset human activity recognition. LanHAR consists of three key components. (i) Generation: we carefully design prompts to guide LLMs in generating semantic interpretations and develop an iterative re-generation method to ensure high-quality semantic outputs. (ii) Alignment: we design a semantic interpretation alignment module to enhance understanding of semantic interpretations based on a text encoder to encode interpretations and two contrastive learning tasks to improve the alignment between activities and sensor readings. (iii) Deployment: We introduce a lightweight, two-stage training and inference framework. First, we train the semantic interpretation alignment module, then design and train a sensor reading encoder to map sensor data to the semantic interpretation space. For inference on mobile devices, we use the trained encoder to obtain sensor reading encodings and measure their similarity to activity label semantics to determine the HAR results.
@@
  • To our knowledge, we are the first to leverage LLMs to generate semantic interpretations for cross-dataset human activity recognition. Our method introduces a novel perspective by utilizing semantic interpretations from both sensor readings and activity labels, which helps mitigate the impact of cross-dataset heterogeneity and opens the door to recognizing new activities.
  • We design LanHAR, a novel system that leverages LLMs to generate semantic interpretations of sensor readings and activity labels, addressing the challenge of cross-dataset human activity recognition. The system features a semantic interpretation generation process with an iterative re-generation method to ensure high-quality outputs, alongside a two-stage training framework that transfers the capabilities of large language models to mobile devices.
  • We evaluate our system based on five public datasets. The experimental results demonstrate that it outperforms other state-of-the-art methods in cross-dataset activity recognition, achieving an improvement of 7.35% in accuracy and 13.16% in F1 score. Additionally, for new activity recognition, it shows further enhancement, with improvements of 43.67% in accuracy.
@@ 2.1. LLMs possess the ability to perceive the physical world
Existing research (Xu et al., 2024; Ji et al., 2024) has demonstrated that LLMs possess the ability to perceive and interpret the physical world. For instance, LLMs can analyze various types of sensor data to infer a person’s location or activity. Building on this capability, we explore the use of LLMs to analyze sensor readings and describe activity labels, thereby generating semantic interpretations for both. Specifically, we design prompts that include data introduction, data analysis, relevant knowledge, and task instructions (detailed in Section 3.3) to guide LLMs in generating semantic interpretations of sensor readings. Similarly, we guide LLMs to describe activity labels. Fig. 2 provides an example where GPT-4 generates its understanding of jogging alongside an analysis of sensor data sequences.
@@ 2.2. Semantic interpretations of sensor readings for cross-dataset HAR
We explain why LLM-generated semantic interpretation offers a valuable method to address data heterogeneity in cross-dataset HAR. The key intuition is that, despite variations in how heterogeneous data manifests, there are often underlying common activity patterns. By identifying these shared patterns, we can map the original heterogeneous sensor data into a common shared space (i.e., a language space with encoded semantic interpretations), effectively reducing the impact of data heterogeneity.
@@
To understand how data heterogeneity is mitigated using generated semantic interpretations, we compare data distributions under three settings: raw data, data encoded by encoders, and generated semantic interpretations. Data encoding is a common method in existing research (Auge et al., 2021), aimed at learning better representations or embeddings of sensor readings. For data encoding, we employ a self-supervised learning model, BERT (Devlin et al., 2018), as a representative method, following prior work (Xu et al., 2021) that masks and predicts parts of the sensor readings. For semantic interpretation, we use GPT-4 to generate interpretations from raw IMU data based on a designed prompt (detailed in Section 3.3). These interpretations are then input into a pre-trained language model (BERT) to produce representations in the language space. We use two public datasets—UCI (Reyes-Ortiz et al., 2016) (D1) and Motion (Malekzadeh et al., 2019) (D2) as a cross-dataset example including four common activities: walking, sitting, going upstairs, and going downstairs.
@@
Qualitative results: We visualize the results from the three settings using t-distributed Stochastic Neighbor Embedding (t-SNE) (Van der Maaten and Hinton, 2008) to project them onto a 2D surface. As an example, we plot two activities, “sit” (blue) and “walk” (orange), in Fig. 3. Initially, in the raw data, the same activities from different datasets are far apart. After applying self-supervised learning, the distribution gap between the datasets decreases, but the activities also become less distinguishable. However, when using semantic interpretations, the distribution gap reduces further, and the same activities (same color) cluster more closely, demonstrating improved grouping of the same activities.
@@
Quantitative results: We quantitatively assess the impact of introducing semantic interpretations for sensor readings across all activities. To measure the distribution differences between the two datasets under the three settings, we use relative Kullback-Leibler (KL) divergence, computed as the ratio of KL divergence between the distributions of the same activity to KL divergence between the distributions of different activities (Kullback and Leibler, 1951), a widely used metric for comparing distributions. Compared to KL divergence, relative KL divergence normalizes differences in scale, mitigates the inherent closeness of word embeddings, and ensures a fair comparison across activities. Table 1 presents the relative KL divergence for all activities across the two datasets under each setting. The results show that introducing semantic interpretations significantly reduces the distribution gap for all activities. On average, the KL divergence is reduced by 40.45%.
@@
Table 1. Relative KL divergence between the distributions of the same activity across two datasets
@@
Walking 0.4965 0.2331 0.1633
Sitting 0.1669 0.1983 0.1541
Going upstairs 0.2444 0.1875 0.0420
Going downstairs 0.1753 0.1518 0.0995
Average 0.2708 0.1926 0.1147
@@ 2.3. New activity recognition
Recognizing new activities has been a longstanding challenge in HAR, particularly for machine learning-based methods, which are unable to identify activities that were not present during the training phase. The most relevant method to address this challenge involves using predefined attributes of activities to establish connections between new and existing activities (Cheng et al., 2013). However, the performance of this method relies heavily on the quality of manually defined activity attributes, making it costly and difficult to generalize to new labels.
We argue that combining the semantic interpretations of sensor readings and labels offers a new perspective for addressing this challenge. By converting all activity labels into a language space using their semantic interpretations, we can represent labels even without corresponding sensor readings (which is common for new labels). Similarly, with sensor readings mapped into the same language space, we can match each reading to a label, regardless of whether the label was previously encountered. Fig. 4 illustrates this concept, where we assume “biking” and “going upstairs” are known activities from the training phase, while “going downstairs” is a new activity. In our method, we first convert all activity labels into the same language space using their LLM-generated semantic interpretations. We then obtain semantic interpretations for the sensor readings. Even though “going downstairs” is unseen during training, we can still measure its distance to all activity labels in the language space, enabling the recognition of new activities. In our setting, we assume that we have the knowledge of possible activity labels in the inference stage.
@@ 3.1. Problem formulation
Figure 5. Overview of LanHAR. (1) Utilize LLMs for generating semantic interpretations of sensor reading and activity labels. (2) Train a text encoder to encode two types of semantic interpretations and achieve their alignment. $H{i}$ and $Z_{i}$ denote embeddings of semantic interpretations of activity labels and sensor reading. (3) Train a sensor encoder to align sensor reading and semantic interpretation. $E{i}$ denote embeddings of sensor reading. (4) For inference, only use the sensor encoder to generate embeddings for the sensor readings $E{i}$ and then compute the similarity with the pre-stored embeddings of the activity labels $H{i}$ to obtain the human activity recognition results.
@@ 3.2. Design overview
We design LanHAR, a novel system that uses LLM-generated semantic interpretations of sensor readings and activity labels for cross-dataset human activity recognition. As shown in Fig. 5, LanHAR consists of four processes:
@@ 3.3. LLMs for semantic interpretations
We introduce how we guide LLMs to generate semantic interpretations of sensor readings and activity labels. We also discuss our method to handling low-quality interpretations to ensure the reliability of LLM responses.
@@ 3.3.1. Obtaining semantics interpretation of sensor reading
We carefully design a prompt to help LLMs better understand the problem setting and deliver more accurate and meaningful answers. An example is shown in Fig. 6. The prompt consists of four key parts: data introduction, data analysis, knowledge, and task introduction.
@@ 3.3.2. Obtaining semantics interpretation of activity label
Similarly, we design a prompt to guide LLMs in generating semantic interpretations of activity labels. In this prompt, we outline the task for the LLMs, instructing them to generate a description for a given activity. The description should cover three key aspects: a general overview of the activity, the potential states or patterns detected by the accelerometer and gyroscope during the activity, and the body parts likely involved in performing the activity. An example is shown in Fig. 7.
@@ 3.3.3. An iterative re-generation method to ensuring the quality of LLM responses
The quality of LLMs’ interpretations is critical for our method. However, due to hallucinations or other incidental factors, LLMs may sometimes generate inaccurate or illogical responses. Through manual review, we identify numerous instances of this issue when generating semantic interpretations. To improve the accuracy of LLM responses, we develop an iterative re-generation method to ensure high-quality outputs. The key intuition is that refining responses through clustering-based filtering will lead to a lower intra-cluster distance within the same activity. As illustrated in Fig. 8, the framework first clusters the responses and identifies the $k$ most problematic ones based on their deviation from the cluster center. These data points are re-input into the LLM to regenerate the semantic interpretations. We then assess whether to incorporate the new interpretations based on its impact on the intra-cluster distance. The details are as follows. We show in the evaluation that this process improves the data quality.
Figure 8. The iterative re-generatio method to ensuring the quality of LLM responses. (1) Filter inaccurate semantic interpretations. (2) Regenerate new semantic interpretations by LLMs. (3) Incorporate the new semantic interpretation based on its impact on the intra-cluster distance.
(1) Filter out inaccurate semantic interpretations: We apply a clustering-based method (i.e., K-means) to identify anomalous data points within the generated semantic interpretations of the same activity. After clustering, we select $k$ outliers based on their distance from the cluster center, with those farther than the average intra-cluster distance considered low-quality responses for refinement.
(2) Regenerate new semantic interpretations by LLMs: After identifying the $k$ outliers, we update the task introduction in the prompt (Section 3.3.1) to inform the LLMs of the inaccurate responses. The revised prompt includes the following: “This is your previous response to the task. Please analyze it step by step to identify any logical errors, inconsistencies with real-world knowledge, or discrepancies with the input data. Provide a corrected response according to the required format.”
(3) Incorporate the new semantic interpretation based on its impact on the cluster structure: If the newly generated interpretations reduce the intra-cluster distance, making the responses within the same cluster more compact, we accept the new semantic interpretation; otherwise, we retain the original one. We repeat this process until either the cluster structure remains stable (e.g., the intra-cluster distance change falls below a predefined threshold) for multiple continuous iterations or the specified number of iterations is reached.
Note that the iterative re-generation process occurs only during the offline training stage and does not impact the efficiency of online inference when deployed on mobile devices. In our study, only 13.30% of the semantic interpretations were identified as inaccurate.